# Distributional Lusternik–Schnirelmann category of manifolds

### v2

@@ 8.B. Using curvature and cohomogeneity
Before we make our observation, we describe some flat $3$-manifolds (v)(v)(v) As in [LS], we use the notations ${\mathcal{M}}_{i}$ and $\iota_{j}$ for manifolds and their involutions, respectively..
Let $S^{1}=\{z\in{\mathbb{C}}\mid|z|=1\}$ be the unit circle and $T^{2}=S^{1}\times S^{1}$ be the $2$-torus. Define ${\mathcal{M}}_{2}=T^{2}\times[-1,1]/\sim$, where $(z_{1},z_{2},1)\sim(\overline{z_{1}},\overline{z_{2}},-1)$. Then ${\mathcal{M}}_{2}$ is a $T^{2}$-bundle over $S^{1}$. It is a closed orientable flat $3$-manifold with holonomy group ${\mathbb{Z}}_{2}$. Let us define $\iota_{4}:{\mathcal{M}}_{2}\to{\mathcal{M}}_{2}$ as $\iota_{4}([z_{1},z_{2},t])=[-\overline{z_{1}},\overline{z_{2}},-t]$. It is an orientation-reversing isometric involution having four isolated fixed points.
Let $W=S^{1}\times[0,1]\times[-1,1]/\sim$, where $(z,0,t)\sim(\overline{z},1,-t)$. So, $W$ is an orientable twisted interval-bundle over a Klein bottle. Let $W_{1}$ and $W_{2}$ be two copies of $W$, and let $g:\partial W_{1}\to\partial W_{2}$ be the homeomorphism $g([z_{1},z_{2},1])=[z_{2},z_{1},1]$. Define ${\mathcal{M}}_{6}=(W_{1}\sqcup W_{2})/\approx$, where $x\approx g(x)$ for all $x\in\partial W_{1}$. Then ${\mathcal{M}}_{6}$ is the closed orientable Hantzsche–Wendt $3$-manifold discussed in Example 4.8. Define a map $\iota_{2}:{\mathcal{M}}_{6}\to{\mathcal{M}}_{6}$ by pasting such that $\iota_{2}([z_{1},z_{2},t])=[-\overline{z_{1}},-z_{2},t]$ on $W_{1}$ and $\iota_{2}([z_{1},z_{2},t])=[-z_{1},-\overline{z_{2}},t]$ on $W_{2}$. Then, $\iota_{2}$ is an orientation-reversing isometric involution having two isolated fixed points.
@@
Since $\operatorname{Susp}({\mathbb{R}}P^{2})$ and $\operatorname{Susp}({\mathbb{R}}P^{2})\#\operatorname{Susp}({\mathbb{R}}P^{2})$ are simply connected CW complexes, their LS-category is $1$. So, by Theorem 8.3, there must exist a closed orientable flat $3$-manifold $M$ and an orientation-reversing isometric involution $\iota_{M}$ with isolated fixed points such that $X=M/\iota_{M}$. Due to Luft and Sjerve [LS], the only possibilities are $M\in\{{\mathcal{M}}_{2},{\mathcal{M}}_{6},T^{3}\}$ and $\iota_{{\mathcal{M}}_{2}}=\iota_{4}$, $\iota_{{\mathcal{M}}_{6}}=\iota_{2}$, or $\iota_{T^{3}}=\iota_{8}$, where $\iota_{8}:T^{3}\to T^{3}$ defined as $\iota_{8}(z_{1},z_{2},z_{3})=(\overline{z_{1}},\overline{z_{2}},\overline{z_{3}})$ has eight fixed points. By [GGG, Page 9], $T^{3}/\iota_{8}$ is a simply connected CW complex. So, $\operatorname{\mathsf{cat}}(T^{3}/\iota_{8})=1$ and thus, $X\neq T^{3}/\iota_{8}$. Hence, either $X={\mathcal{M}}_{2}/\iota_{4}$ or $X={\mathcal{M}}_{6}/\iota_{2}$. ∎
@@
For a closed Alexandrov $3$-space $X$, let $\text{Iso}(X)$ denote the compact Lie group of isometries of $X$. Then the dimension of the orbit space $X/\text{Iso}(X)$ is called the cohomogeneity of the action. We denote it by $c(X)$. Let $G$ be a compact Lie group acting isometrically on $X$. Then $c(X)\in\{0,1,2,3\}$.
@@
Let $X$ be a closed Alexandrov $3$-space with an effective and isometric action of the circle group $S^{1}$. Then $X$ has $2r$ number of singular points for some $r\geq 0$ and $X$ is weakly equivariantly homeomorphic to #$\#_{i=1}^{r}\operatorname{Susp}({\mathbb{R}}P^{2})\#\mathscr{M}$, where $\mathscr{M}$ is the closed $3$-manifold described by the set of invariants $(b;(\epsilon,g,f+s,t);\{(\alpha_{i},\beta_{i})\}_{i=1}^{n})$.
@@
We now give two instances where the converse of Proposition 8.7 holds true.
@@
Let $X=\#_{i=1}^{r}\operatorname{Susp}({\mathbb{R}}P^{2})\#\mathscr{M}$ for some $r\geq 1$ as in Theorem 8.6. If $\mathscr{M}$ is aspherical, then we have $\operatorname{\mathsf{dcat}}(X)=\operatorname{\mathsf{cat}}(X)=3$ by Theorem 7.3. Next, let $X=\operatorname{Susp}({\mathbb{R}}P^{2})\#\operatorname{Susp}({\mathbb{R}}P^{2})\#(S^{2}\times S^{1})$ with one of the two effective and isometric $S^{1}$ actions described in [NZ, Example 5.3]. It follows from the technique of the proof of Lemma 8.1 that $2=c\ell_{{\mathbb{Q}}}(S^{2}\times S^{1})\leq\operatorname{\mathsf{dcat}}(X)\leq\operatorname{\mathsf{cat}}(X)$.
9. Small values of distributional topological complexity
The notion of the distributional topological complexity $(\operatorname{\mathsf{dTC}})$ of a space was introduced in [DJ] (and independently in [KW] as “analog topological complexity”) and studied recently in [J1] and [Dr2]. In [DJ], $\operatorname{\mathsf{dTC}}$ was defined as the distributional version of Farber’s classical notion of the topological complexity of a space [Far] with motivations similar to those for defining $\operatorname{\mathsf{dcat}}$ — see Section 2.A. In this section, we explore some necessary conditions for the $\operatorname{\mathsf{dTC}}$ of a closed orientable manifold to be small, i.e., at most $3$.
Let ${\mathcal{P}}(X)$ denote the full path space of $X$ with the compact-open topology and for any pair $(x,y)\in X\times X$, let ${\mathcal{P}}(x,y)=\{\phi\in{\mathcal{P}}(X)\mid\phi(0)=x,\phi(1)=y\}$. We refer to [DJ, Sections 3.3 and 4.3] for the following discussion.
9.1 Definition.
The distributional topological complexity of a space $X$, denoted $\operatorname{\mathsf{dTC}}(X)$, is the smallest integer $n$ for which there exists a continuous mapping $s:X\times X\to{\mathcal{B}}_{n+1}({\mathcal{P}}(X))$ such that $s(x,y)\in{\mathcal{B}}_{n+1}({\mathcal{P}}(x,y))$ for all $(x,y)\in X\times X$.
Given a space $X$, $\operatorname{\mathsf{dcat}}(X)\leq\operatorname{\mathsf{dTC}}(X)\leq\operatorname{\mathsf{TC}}(X)$, where $\operatorname{\mathsf{TC}}(X)$ denotes the topological complexity of $X$ introduced by Farber in [Far]. The last inequality can be strict. Indeed, $\operatorname{\mathsf{dTC}}({\mathbb{R}}P^{n})=1<n\leq\operatorname{\mathsf{TC}}({\mathbb{R}}P^{n})$ for each $n\geq 2$, in stark contrast to [FTY, Theorem 12]. Given the diagonal map $\Delta:X\to X\times X$ and any ring $R$, the cup length of the ideal $\text{Ker}(\Delta^{*}:H^{*}(X\times X;R)\to H^{*}(X;R))$ is called the $R$-zero-divisor cup length of $X$ and is denoted by $zc\ell_{R}(X)$. While we have $zc\ell_{R}(X)\leq\operatorname{\mathsf{TC}}(X)$ for all rings $R$, in the case of $\operatorname{\mathsf{dTC}}$, we only have that $zc\ell_{\mathbb{F}}(X)\leq\operatorname{\mathsf{dTC}}(X)$ for $\mathbb{F}\in\{{\mathbb{Q}},{\mathbb{R}}\}$.
We note that like $\operatorname{\mathsf{dcat}}$, the (sequential) distributional topological complexity of a space is also a special case of the generalized notion of the “distributional sectional category” of a fibration, see [J1, Section 5]. We refer the reader to [DJ] and [J1] for more details and results on (sequential) distributional topological complexity, see also [KW] for (sequential) analog topological complexity and “analog sectional category”, which are notions closely related to the respective distributional notions, [KW, Conjecture 1.2][J1, Section 8.A], and [Dr2, Page 2].
If $X$ is a finite CW complex, then $\operatorname{\mathsf{TC}}(X)=1\implies X\simeq S^{2k-1}$ for some $k\geq 1$ due to [GLO]. In view of the equality $\operatorname{\mathsf{dTC}}({\mathbb{R}}P^{n})=1$ for all $n\geq 1$, this conclusion does not hold in the theory of $\operatorname{\mathsf{dTC}}$. So, a natural next step should be to find the conclusions that can be derived for a closed manifold $M$ if $\operatorname{\mathsf{dTC}}(M)$ is small but not necessarily $1$. This section is aimed in that direction and is motivated by the recent work of Pavešić in the theory of classical topological complexity. In [Pa], Pavešić gave a complete description of the integral cohomology ring of closed orientable manifolds $M$ for which $\operatorname{\mathsf{TC}}(M)\leq 2$. It turns out that many of his conclusions from the classical theory (vi)(vi)(vi) We note that here, we are working with normalized $\operatorname{\mathsf{dTC}}$ and $\operatorname{\mathsf{TC}}$ which are 00 if and only if the space is contractible, [DJ][GLO]. So, our $\operatorname{\mathsf{TC}}$ is one less than the $\operatorname{\mathsf{TC}}$ studied in [Far] and [Pa]. remain true in the theory of $\operatorname{\mathsf{dTC}}$ as well.
Let $M$ denote a closed orientable manifold of dimension $n\geq 1$ such that $\operatorname{\mathsf{dTC}}(M)\leq 3$. If $n=1$, then $M=S^{1}$, and if $n=2$, then $M=S^{2}$ or $M=S^{1}\times S^{1}$ since $\operatorname{\mathsf{dTC}}$ of each closed orientable surface is known, [DJ, Proposition 6.9]. The rational cohomology rings and fundamental groups of these spaces are well-known. Therefore, we now focus on dimensions $n\geq 3$.
For any $x\in H^{*}(M;{\mathbb{Q}})$, we let $\widehat{x}=x\otimes 1-1\otimes x\in H^{*}(M\times M;{\mathbb{Q}})$ denote the rational zero-divisor corresponding to $x$.
9.2 Proposition.
If $\operatorname{\mathsf{dTC}}(M)\leq 3$, then $\operatorname{rank}(H_{1}(M;{\mathbb{Q}}))\leq 1$.
Proof.
Since $M$ is orientable, by [Ha, Proposition 3.38], the cup product pairing
$H^{1}(M;{\mathbb{Q}})\times H^{n-1}(M;{\mathbb{Q}})\to H^{n}(M;{\mathbb{Q}})$
is non-singular. If $\operatorname{rank}(H_{1}(M;{\mathbb{Q}}))\geq 2$, then by Poincaré duality and the non-singularity of the above pair, there exists $a,b\in H^{1}(M;{\mathbb{Q}})$ and $u,v\in H^{n-1}(M;{\mathbb{Q}})$ such that $a\smile u=b\smile v\neq 0$ and $a\smile v=0=b\smile u$. Then,
$\widehat{a}\smile\widehat{u}\smile\widehat{b}\smile\widehat{v}=2\left((a\smile u)\otimes(b\smile v)\right)\neq 0.$
Thus, we get $4\leq zc\ell_{{\mathbb{Q}}}(M)\leq\operatorname{\mathsf{dTC}}(M)$. This contradicts our hypothesis that $\operatorname{\mathsf{dTC}}(M)\leq 3$. Therefore, $\operatorname{rank}(H_{1}(M;{\mathbb{Q}})\leq 1$. ∎
More can be said if $\operatorname{\mathsf{dTC}}(M)$ values are even smaller.
9.3 Corollary.
If $\operatorname{\mathsf{dTC}}(M)\leq 2$ and $\dim(M)=n$ is odd, then $\operatorname{rank}(H_{1}(M;{\mathbb{Q}}))=0$.
Proof.
Due to Proposition 9.2, $\operatorname{rank}(H_{1}(M;{\mathbb{Q}}))\leq 1$. If $\operatorname{rank}(H_{1}(M;{\mathbb{Q}}))=1$, then from the above non-singular pairing, we have $a\in H^{1}(M;{\mathbb{Q}})$ and $u\in H^{n-1}(M;{\mathbb{Q}})$ such that $a\smile u\neq 0$. Since $n-1$ is even,
$\widehat{a}\smile\widehat{u}\smile\widehat{u}=2\left(a\otimes(a\smile u)-(a\smile u)\otimes a\right)\neq 0.$
Thus, we get $3\leq zc\ell_{{\mathbb{Q}}}(M)\leq\operatorname{\mathsf{dTC}}(M)$. But this contradicts our hypothesis that $\operatorname{\mathsf{dTC}}(M)\leq 2$. Hence, $\operatorname{rank}(H_{1}(M;{\mathbb{Q}}))=0$. ∎
9.4 Corollary.
If $\operatorname{\mathsf{dTC}}(M)\leq 2$, $\dim(M)=n$ is even, and $H^{i}(M;{\mathbb{Q}})\neq 0$ for some $i\in\{2,\ldots,n-2\}$, then $\operatorname{rank}(H_{1}(M;{\mathbb{Q}}))=0$.
Proof.
Due to Proposition 9.2, $\operatorname{rank}(H_{1}(M;{\mathbb{Q}}))\leq 1$. By our assumption, there exists $0\neq w\in H^{i}(M;{\mathbb{Q}})$. If $\operatorname{rank}(H_{1}(M;{\mathbb{Q}}))=1$, then since $n-1$ is odd, $\widehat{a}\smile\widehat{u}\smile\widehat{w}\neq 0$ for the cohomology classes $a\in H^{1}(M;{\mathbb{Q}})$ and $u\in H^{n-1}(M;{\mathbb{Q}})$ as above, see the proof of [Pa, Theorem 3.2]. Thus, $3\leq zc\ell_{{\mathbb{Q}}}(M)\leq\operatorname{\mathsf{dTC}}(M)$. As before, this is a contradiction. Hence, $\operatorname{rank}(H_{1}(M;{\mathbb{Q}}))=0$. ∎
9.5 Remark.
If $\pi:=\pi_{1}(M)$ is free, then all the above conclusions hold for $\operatorname{rank}(\pi)$ as well since in that case, $\operatorname{rank}(\pi)\leq\operatorname{rank}(H_{1}(M;{\mathbb{Q}}))$. In the case of classical $\operatorname{\mathsf{TC}}$, it follows from [DKR, Theorem 4.1] that $\operatorname{\mathsf{TC}}(M)\leq 2\implies\pi$ is free. However, since $\operatorname{\mathsf{dTC}}({\mathbb{R}}P^{n})=1$ for each $n$, such a conclusion does not hold in the theory of $\operatorname{\mathsf{dTC}}$.
Just like in [Pa, Theorem 3.2], here also the structure of the rational cohomology ring of $M$, denoted $H^{*}(M;{\mathbb{Q}})$, can be determined when the $\operatorname{\mathsf{dTC}}(M)$ values are small. However, we will need some additional assumptions. For any ring $R$, let
$\widetilde{M}(R):=\bigoplus_{i=2}^{n-2}H_{i}(M;R).$
If $y\in H^{i}(M;{\mathbb{Q}})$, we write $y_{i}$ to denote that $y$ is a cohomology class of degree $i$.
9.6 Theorem.
Let $M$ be a closed orientable $n$-manifold such that $\operatorname{\mathsf{dTC}}(M)\leq 2$. If $\pi:=\pi_{1}(M)$ is free and $\widetilde{M}({\mathbb{Z}}_{p})=0$ for all primes $p\geq 2$, then either
  1. (1)
    $H^{*}(M;{\mathbb{Q}})=\wedge_{{\mathbb{Q}}}(x_{n})$, or
  2. (2)
    $H^{*}(M;{\mathbb{Q}})=\wedge_{{\mathbb{Q}}}(x_{k},x_{l})$ for $k,l$ odd such that $l\geq 3$ and $k+l=n$.
Proof.
Since $\operatorname{\mathsf{dTC}}(M)\leq 2$ and $\pi$ is free, $\operatorname{rank}(\pi)\in\{0,1\}$ by Proposition 9.2 and Remark 9.5. So, the only possibilities are $\pi=0$ or $\pi={\mathbb{Z}}$. First, assume $\pi={\mathbb{Z}}$. Then, $\dim(M)=n$ is even and $\widetilde{M}({\mathbb{Q}})=0$ due to Corollaries 9.3 and 9.4, respectively. Hence, as noted in the proof of [Pa, Theorem 3.2], $H^{*}(M;{\mathbb{Q}})=\wedge_{{\mathbb{Q}}}(x_{1},x_{n-1})$ where $x_{1}$ and $x_{n-1}$ are Poincaré duals to each other. The only remaining case is when $\pi=0$. In this case, $H_{1}(M;R)=0=H_{n-1}(M;R)$ for each coefficient ring $R$ due to the universal coefficient theorems and Poincaré duality. We now have the following two situations.
  • If $\widetilde{M}({\mathbb{Q}})=0$, then $H^{*}(M;{\mathbb{Q}})=\wedge_{{\mathbb{Q}}}(x_{n})$ since $\widetilde{M}({\mathbb{Z}}_{p})=0$ for all primes $p\geq 2$ by our assumption.
  • If $\widetilde{M}({\mathbb{Q}})\neq 0$, then due to [Ha, Corollary 3.39], there exist cohomology classes $a,b\in H^{*}(M;{\mathbb{Q}})$ of infinite order such that $a\smile b\neq 0$. If the degree of either $a$ or $b$ is even, then we can proceed as in the proof of Corollary 9.3 to get $\operatorname{\mathsf{dTC}}(M)\geq 3$. This contradicts our hypothesis that $\operatorname{\mathsf{dTC}}(M)\leq 2$. Thus, both $a$ and $b$ must have odd degrees. Then as noted in the proof of [Pa, Theorem 3.2], the ring $\wedge_{{\mathbb{Q}}}(x_{k},x_{l})$ is a subring of $H^{*}(M;{\mathbb{Q}})$ for $k,l$ odd such that $3\leq k\leq l<n$ and $k+l=n$. In fact, we must have $H^{*}(M;{\mathbb{Q}})=\wedge_{{\mathbb{Q}}}(x_{k},x_{l})$, for if there exists $c\in H^{*}(M;{\mathbb{Q}})\setminus\wedge_{{\mathbb{Q}}}(x_{k},x_{l})$ , then $\widehat{a}\smile\widehat{b}\smile\widehat{c}\neq 0$ as in the proof of Corollary 9.4, contradicting $\operatorname{\mathsf{dTC}}(M)\leq 2$.
Since we have considered all the possibilities, the conclusions follow. ∎
We note that the above characterization of the rational cohomology of $M$ is quite restrictive compared to the characterization of [Pa] because of our assumption $\widetilde{M}({\mathbb{Z}}_{p})=0$ for all primes $p\geq 2$. If $\widetilde{M}({\mathbb{Z}}_{p})\neq 0$ for any prime $p\geq 3$, then $3\leq zc\ell_{{\mathbb{Z}}_{p}}(M)\leq\operatorname{\mathsf{TC}}(M)$. However, since $zc\ell_{{\mathbb{Z}}_{p}}(M)$ is not necessarily a lower bound to $\operatorname{\mathsf{dTC}}(M)$, the possibility $\widetilde{M}({\mathbb{Z}}_{p})\neq 0$ for odd primes $p$ cannot be eliminated here.

### v3

@@ 8.2. Using curvature and cohomogeneity
Before we make our observation, we describe some flat $3$-manifolds (i)(i)(i) As in [LS], we use the notations ${\mathcal{M}}_{i}$ and $\iota_{j}$ for manifolds and their involutions, respectively..
Let $S^{1}=\{z\in{\mathbb{C}}\mid|z|=1\}$ be the unit circle and $T^{2}=S^{1}\times S^{1}$ be the $2$-torus. Define ${\mathcal{M}}_{2}=T^{2}\times[-1,1]/\sim$, where $(z_{1},z_{2},1)\sim(\overline{z_{1}},\overline{z_{2}},-1)$. Then ${\mathcal{M}}_{2}$ is a $T^{2}$-bundle over $S^{1}$. It is a closed orientable flat $3$-manifold with holonomy group ${\mathbb{Z}}_{2}$. Let us define $\iota_{4}:{\mathcal{M}}_{2}\to{\mathcal{M}}_{2}$ as
$\iota_{4}([z_{1},z_{2},t])=[-\overline{z_{1}},\overline{z_{2}},-t].$
It is an orientation-reversing isometric involution having four isolated fixed points.
Let $W=S^{1}\times[0,1]\times[-1,1]/\sim$, where $(z,0,t)\sim(\overline{z},1,-t)$. So, $W$ is an orientable twisted interval-bundle over a Klein bottle. Let $W_{1}$ and $W_{2}$ be two copies of $W$, and let $g:\partial W_{1}\to\partial W_{2}$ be the homeomorphism $g([z_{1},z_{2},1])=[z_{2},z_{1},1]$. Define ${\mathcal{M}}_{6}=(W_{1}\sqcup W_{2})/\approx$, where $x\approx g(x)$ for all $x\in\partial W_{1}$. Then ${\mathcal{M}}_{6}$ is the closed orientable Hantzsche–Wendt $3$-manifold discussed in Example 4.8. Define a map $\iota_{2}:{\mathcal{M}}_{6}\to{\mathcal{M}}_{6}$ such that
$\iota_{2}([z_{1},z_{2},t])=\begin{cases}[-\overline{z_{1}},-z_{2},t]&\text{ if}[z_{1},z_{2},t]\in W_{1}\\ [-z_{1},-\overline{z_{2}},t]&\text{ if }[z_{1},z_{2},t]\in W_{2}.\end{cases}$
Then, $\iota_{2}$ is an orientation-reversing isometric involution having two isolated fixed points.
@@
Since $\operatorname{Susp}({\mathbb{R}}P^{2})$ and $\operatorname{Susp}({\mathbb{R}}P^{2})\#\operatorname{Susp}({\mathbb{R}}P^{2})$ are simply connected CW complexes, their LS-category is $1$. So, by Theorem 8.3, there must exist a closed orientable flat $3$-manifold $M$ and an orientation-reversing isometric involution $\iota_{M}$ with isolated fixed points such that $X=M/\iota_{M}$. Due to Luft and Sjerve [LS], the only possibilities are $M\in\{{\mathcal{M}}_{2},{\mathcal{M}}_{6},T^{3}\}$ and $\iota_{{\mathcal{M}}_{2}}=\iota_{4}$, $\iota_{{\mathcal{M}}_{6}}=\iota_{2}$, or $\iota_{T^{3}}=\iota_{8}$, where $\iota_{8}:T^{3}\to T^{3}$ defined as the involution
$\iota_{8}(z_{1},z_{2},z_{3})=(\overline{z_{1}},\overline{z_{2}},\overline{z_{3}})$
which has eight fixed points. By [GGG, Page 9], $T^{3}/\iota_{8}$ is a simply connected CW complex. Hence, $\operatorname{\mathsf{cat}}(T^{3}/\iota_{8})=1$ and thus, $X\neq T^{3}/\iota_{8}$. Therefore, we conclude that either $X={\mathcal{M}}_{2}/\iota_{4}$ or $X={\mathcal{M}}_{6}/\iota_{2}$. ∎
@@
For a closed Alexandrov $3$-space $X$, let $\text{Iso}(X)$ denote the compact Lie group of isometries of $X$. Then the dimension of the orbit space $X/\text{Iso}(X)$ is called the cohomogeneity of the action. We denote it by $c(X)$. Clearly, $c(X)\in\{0,1,2,3\}$. Let $G$ be a compact Lie group acting isometrically on $X$. If $G\neq 0$ acts effectively on $X$, then $c(X)\neq 3$.
@@
Let $X$ be a closed Alexandrov $3$-space with an effective and isometric action of the circle group $S^{1}$. Then $X$ has $2r$ number of singular points for some $r\geq 0$ and $X$ is weakly equivariantly homeomorphic to
#$\#_{i=1}^{r}\operatorname{Susp}({\mathbb{R}}P^{2})\#\mathscr{M},$
where $\mathscr{M}$ is the closed manifold of dimension $3$ described by the set of invariants $(b;(\epsilon,g,f+s,t);\{(\alpha_{i},\beta_{i})\}_{i=1}^{n})$.
@@
We end this paper by giving two instances where the converse of Proposition 8.7 holds true.
@@
Let $X=\#_{i=1}^{r}\operatorname{Susp}({\mathbb{R}}P^{2})\#\mathscr{M}$ for some $r\geq 1$ as in Theorem 8.6. If $\mathscr{M}$ is aspherical, then we have $\operatorname{\mathsf{dcat}}(X)=\operatorname{\mathsf{cat}}(X)=3$ by Theorem F. Next, let
$X=\operatorname{Susp}({\mathbb{R}}P^{2})\#\operatorname{Susp}({\mathbb{R}}P^{2})\#(S^{2}\times S^{1})$
with any one of the two effective and isometric $S^{1}$ actions described explicitly in [NZ, Example 5.3]. It follows from the technique of the proof of Lemma 8.1 that $2=c\ell_{{\mathbb{Q}}}(S^{2}\times S^{1})\leq\operatorname{\mathsf{dcat}}(X)\leq\operatorname{\mathsf{cat}}(X)$.